# From Calculation to Adjudication: Examining LLM judges on Mathematical Reasoning Tasks

### v1

@@ Abstract
To reduce the need for human annotations, large language models (LLMs) have been proposed as judges of the quality of other candidate models. LLM judges are typically evaluated by measuring the correlation with human judgments on generation tasks such as summarization or machine translation. In contrast, we study LLM judges on mathematical reasoning tasks. These tasks require multi-step reasoning, and the correctness of their solutions is verifiable, enabling a more objective evaluation. We perform a detailed performance analysis and find that the used judges are mostly unable to improve task performance but are able to pick the better model. Our analysis uncovers a strong correlation between judgment performance and the candidate model task performance. We observe that judges tend to choose the model of higher quality even if its answer is incorrect. Further, we show that it is possible to use statistics, such as the task performances of the individual models, to predict judgment performance. In an ablation, we either swap or mask the candidate answers and observe that judges often keep the original judgment, providing evidence that judges incorporate writing style in their judgments. In summary, we find that regularities in the judgments are quantifiable using statistical measures and provide various angles on exploiting them. [^1] Code: git@github.com:AndSt/llm_judges.git
@@ 1 Introduction
The automatic evaluation of machine learning models promises to reduce the need for human annotations. Specifically, the LLM-as-a-judge paradigm Zheng et al. (2023) has gained traction, aiming to assess or compare the quality of generated texts automatically. This approach is beneficial for automated data labeling Tan et al. (2024), self-improvement of LLMs Wu et al. (2024), and ranking LLMs with respect to specific tasks Zheng et al. (2023).
@@
Figure 1: In our problem setup two LLMs ($M_{A}$ and $M_{B}$), provide candidate answers for a math problem, and a judge LLM has to decide which one is correct. All three use chain-of-thought (CoT) reasoning Wei et al. (2022).
@@
Much like judges in the real world, who are expected to be exact, fair, and unbiased, e.g., as defined in Bangalore Principles of judicial conduct Bangalore Principles, (2002), LLMs, when employed as judges, should be ethical and logical. Already the philosopher Aristotle argued that the virtuous actor exhibits the joint excellence of reason and character Kraut (2022). Previous works investigate properties and biases of LLM judges on generation tasks such as translation or summarization Kim et al. (2024b); Liu et al. (2024). These are typically evaluated using correlation with human annotators and are thus inherently subjective.
@@
In this work, we investigate LLM judges on mathematical reasoning datasets (see Figure 1). These need complex multi-step reasoning, and the solution is verifiable, which allows us to investigate the relationship between judge and candidate models in a principled manner. We base our analysis on four large (more than 30B parameters) LLMs and four small (less than 10B) LLMs on three mathematical reasoning datasets.
Our experiments contain a detailed performance examination, confirming that larger models are generally better judges Zheng et al. (2023). We find that only the best-tested model, Qwen 2 72B, consistently improves task performance if we evaluate the judged samples, but all tested judges likely pick the better model for a given task.
We investigate subsets with one correct and one incorrect candidate answer. We uncover a correlation between judgment performance and task performance of the candidate models, showing that judges tend to select incorrect answers from better models. Thus, we hypothesize that judges have access and rely on the superior writing styles of larger models instead of solely analyzing the reasoning. When we divide the datasets into buckets of model agreement, we observe that agreement is a proxy for sample difficulty.
Motivated by these regularities, we analyze whether it is possible to predict judgment performance and find that task performances of judge and candidate LLMs explain most of the variance. We hypothesize that judges incorporate writing style into their judgments. Thus, we predict individual judgments using statistical and transformer-based models and achieve above-chance performance, supporting our hypothesis.
Lastly, we test how judgments are affected by perturbing numeric values in responses by 1) swapping results and 2) masking numeric values. Our findings reveal that judges largely retain original judgments, providing further evidence that judges, in large part, base their decisions on writing style. In summary, our contributions are as follows:
@@
    We conduct an in-depth performance analysis of LLM judges for mathematical reasoning tasks.
@@
    Our analysis reveals a correlation between the judgment and candidate task performance, providing a novel statistical angle on the analysis of LLM judges.
@@
    We show that statistics such as task performance or agreement of candidate models are indicative of judgment performance.
@@
    After systematically perturbing the candidate answers, we observe that judges often keep their original judgments, providing evidence that judgments are also based on writing style.
@@ 2.1 LLM as Judges
Using LLMs as judges to evaluate text generated by LLMs, including their own outputs, has recently attracted significant interest because it reduces the need for human annotation Zheng et al. (2023). Commonly, large frontier models are used as judges. Applications include the automatic assessment of language model capabilities and, e.g., determining which model performs better on a given task Zheng et al. (2023) and reinforcement learning from AI feedback by automatically generating data for preference optimization Bai et al. (2022); Wu et al. (2024).
Various methods exist to make judgments Zheng et al. (2023); Liusie et al. (2024). One approach is pairwise selection Wang et al. (2024a), where two answers are presented, and the model is asked to select the better one. Another approach is pointwise grading Li et al. (2024), where the model is asked to assign a grade based on a predefined scale, and the answer with a better grade is chosen. Judgment prompts may involve reference solutions or not. Another body of research explicitly trains models to act as judges Kim et al. (2024a); Wang et al. (2024a) or closely related, as reward models Wang et al. (2024b); Li et al. (2024).
The effectiveness of LLMs as judges is typically assessed by measuring the correlation or overlap with human judgments Zheng et al. (2023); Kim et al. (2024b). In contrast, we focus on difficult tasks with a concrete final answer. Finally, we want to stress that several works caution for the use of LLM judges as experts Bavaresco et al. (2024); Koo et al. (2023); Raina et al. (2024). In a similar vein, we aim to understand regularities and their shortcomings.
@@ 2.2 Biases in LLM-as-a-judge
One linguistic bias is ordering bias Zheng et al. (2023); Koo et al. (2023), where a judge gives a different answer depending on the order in which answers are presented. Panickssery et al. (2024) note that it is possible to interpret position bias as a sign that the model is unsure. There are multiple works Xu et al. (2024); Panickssery et al. (2024); Liu et al. (2024) that find evidence for self-bias or self-preference. Koo et al. (2023) provide a benchmark for analyzing cognitive biases. West et al. (2024) and Oh et al. (2024) explore the “Generative AI Paradox” where generating solutions is easier for the LLM than analyzing them, unlike humans who typically find analysis easier than generation.
In this work, we aim to establish a better understanding of underlying regularities that relate judgments to statistics such as model performance.
@@ 3.1 Problem Description
In this work, we consider two models, denoted by $M_{A},M_{B}\in\mathbf{M}$, providing candidate solutions for a sample of a dataset $D$ and a judge model $M_{J}\in\mathbf{M}$, which is tasked to select, to “judge”, whether it prefers the solutions of the models $M_{A}$ or $M_{B}$. The solutions are represented by the random variables $A$ and $B$. We consider the events that solutions are true $(A=T)$, false $(A=F)$, or that their solution is the same $(A=B)$. We denote the judgment of the judge $M_{J}$ by the random variable $\Delta_{J}$, which can either be correct $(\Delta_{J}=T)$, incorrect $(\Delta_{J}=F)$ or choose a specific model $M_{A}$ $(\Delta_{J}=M_{A})$.
Given that the final answer is either correct or incorrect, we can break the probability of the judge making a correct judgment $P(\Delta_{J}=T|A,B,D)$ given a sample of a dataset $D$ and the answers of two models $M_{A},M_{B}$ down into the following four cases:
$P(\Delta_{J}=T|A,B,D)$ (1)
$=\sum_{(X,Y)\in C}P(\Delta_{J}=T|A=X,B=Y,D)P(A=X,B=Y|D)$
$=P(A=T,B=T|D)$
$+P(\Delta_{J}=T|A=T,B=F,D)P(A=T,B=F|D)$
$+P(\Delta_{J}=T|A=F,B=T,D)P(A=T,B=F|D)$
where $C=(T,F)^{2}$. Note that in cases where both answers are correct or incorrect imply that the judgment is also either correct or incorrect respectively, i.e., $P(\Delta_{J}=T|A=T,B=T)=1$ and $P(\Delta_{J}=T|A=F,B=F)=0$.
@@ 3.3 Models
We evaluate the performance of openly available LLMs, including four large models Qwen 2 72B Yang et al. (2024), Llama 3 70B AI@Meta (2024), Yi 1.5 34B Young et al. (2024), Mixtral 8x7B Jiang et al. (2024) and four small models, namely Llama 3 8B AI@Meta (2024), Gemma 1.1 7B Gemma Team et al. (2024), Mistral 7B v0.3 Jiang et al. (2023), and Mistral 7B v0.1 Jiang et al. (2023). We use the chat- or instruction-tuned model variants and test each model as a candidate answer generator and as a judge. More information is in Appendix A.2.
3.4 Inferences
This section describes the candidate answer generation and the judgment generation. Find more information on prompts and hardware details in Appendix A.
@@ 3.4 Inferences
To judge two candidate answers (including of the same model), we sample two initial CoT solutions for each model using 4-shot prompting. We set the temperature to $0.9$ to get two different solutions.
Qwen 2 72B Llama 3 70B Yi 1.5 34B Mixtral 8x7B Llama 3 8B Gemma 1.1 7B Mistral 7B v0.3 Mistral 7B v0.1
(1) $P(\Delta_{J}=T|A,B,D)$ AQUA_RAT 66.05 55.98 62.2 57.08 51.59 53.52 54.96 51.41
GSM8K 77.06 72.41 72.64 68.81 65.88 65.47 68.57 63.28
MATH 29.66 24.64 26.69 23.60 22.35 21.66 22.68 19.91
(2) $P(\Delta_{J}=T|A\neq B,D)$ AQUA_RAT 53.79 45.04 49.73 44.47 36.56 40.48 38.26 36.43
GSM8K 63.16 59.33 56.01 46.68 41.29 38.71 42.59 39.72
MATH 25.04 22.14 22.22 18.29 17.04 16.00 16.88 15.22
(3) $P(\Delta_{J}=T|\{A,B\}=\{T,F\},D)$ AQUA_RAT 73.13 64.63 68.66 63.26 52.74 58.04 54.54 52.45
GSM8K 85.65 81.32 76.84 64.97 57.87 54.16 59.58 55.48
MATH 80.48 73.01 71.37 61.03 58.04 54.70 56.27 50.91
Table 1: Performance of judge LLMs in three cases: (1) accuracy on all samples, (2) accuracy where models $M_{A}$ and $M_{B}$ disagree, and (3) accuracy where only one model is correct. Results are averaged over all pairs $(M_{A},M_{B})$, with the highest accuracy in bold and the second highest underlined.
@@
We choose the first candidate generation for each model and generate judgments for all $36$ unique model combinations. [^2] We consider all pairs from the eight LLMs, including self-pairing, yielding $\binom{8+2-1}{2}=36$ combinations. If both models are the same, we take the second initial generation. We accommodate positional bias Zheng et al. (2023); Koo et al. (2023) by evaluating the two candidate answers in both possible orders for each question and then taking the average correctness of the judgments as the final assessment. The judge has to choose if the first or second answer is correct. The prompt is zero-shot and applies CoT, the temperature is set to 0 for deterministic generation results.
4 General Performance
The experiments have multiple degrees of freedom: judges, candidate models, and datasets. Therefore, we first examine judgments per dataset, and secondly, we investigate judgments per candidate model pair. Afterwards, we provide evaluations for two applied questions.
4.1 Performance per dataset
We begin by examining the judgment performance, i.e., how often the judge picks a correct answer, across different datasets. Therefore, we average the performance across all model pairs $(M_{A},M_{B})$.
@@ 4.1 Performance per dataset
Table 1 considers three cases where each case focuses on a specific subset of the datasets: Case (1) investigates the observed task performance $P(\Delta_{J}=T|A,B,D)$ where we evaluate the task performance using the answers chosen by the judges. Note that this includes samples where both candidate models give the same answer. Case (2) asks how often judges choose a correct answer given that the answers differ, i.e., $P(\Delta_{J}=T|A\neq B,D)$. Note that this may (and often does) include cases where both answers are incorrect. Case (3) gives the probability that the judge chooses the correct answer given that one answer is correct, and the other answer is incorrect, formally $P(\Delta_{J}=T|A\neq B,T\in\{A,B\},D)$.
@@ Results.
We observe that large models outperform smaller models. Specifically, we see that Qwen 2 72B is the best judge, followed by Yi 1.5 34B. The performance of Llama 3 70B is, on average, comparable to that of Yi 1.5 34B. Note that performance in Case (1) and Case (2) is often quite low, especially for MATH, as there are many cases where the judge can only choose wrong answers. Importantly, we observe that smaller models with fewer than 10B parameters are unreliable judges. Especially, in Case (3), where a correct answer is provided, smaller models only achieve an accuracy of around 55%, barely better than random chance. Therefore, we focus on the four larger models as judges in the subsequent analysis.
(a) Qwen 2 70B
(b) LLama3 70B
(c) Yi 1.5 34B
(d) Mixtral 8x7B
Figure 2: Observed performance $P(\Delta_{J}=T|A,B,D)$ of four judge LLMs (a-d) in evaluating various model pairs, averaged across all datasets.

### v2

@@ Abstract
To reduce the need for human annotations, large language models (LLMs) have been proposed as judges of the quality of other candidate models. The performance of LLM judges is typically evaluated by measuring the correlation with human judgments on generative tasks such as summarization or machine translation. In contrast, we study LLM judges on mathematical reasoning tasks. These tasks require multi-step reasoning, and the correctness of their solutions is verifiable, enabling a more objective evaluation. We perform a detailed performance analysis and find that easy samples are easy to judge, and difficult samples are difficult to judge. Our analysis uncovers a strong correlation between judgment performance and the candidate model task performance, indicating that judges tend to favor higher-quality models even if their answer is incorrect. As a consequence, we test whether we can predict the behavior of LLM judges using simple features such as part-of-speech tags and find that we can correctly predict 70%-75% of judgments. We conclude this study by analyzing practical use cases, showing that LLM judges consistently detect the on-average better model but largely fail if we use them to improve task performance.  [^1] Code will be made available upon acceptance.
@@ 1 Introduction
The automatic evaluation of machine learning models promises to reduce the need for human annotations. Specifically, the LLM-as-a-judge paradigm Zheng et al. (2023) has gained traction, aiming to assess or compare the quality of generated texts automatically. This approach is beneficial for automated data labeling Tan et al. (2024), self-improvement of LLMs Wu et al. (2024), and ranking the capabilities of LLMs, potentially concerning specific tasks Zheng et al. (2023).
@@
Figure 1: In our problem setup two LLMs ($A$ and $B$), provide candidate answers for a math problem, and a judge LLM has to decide which one is correct. All three use chain-of-thought (CoT) reasoning Wei et al. (2022).
@@
Much like judges in the real world, who are expected to be exact, fair, and unbiased Bangalore Principles, (2002), LLM judges, should be unbiased and logical. Previous works investigate properties and biases of LLM judges on generative tasks such as translation or summarization Kim et al. (2024b); Liu et al. (2024), typically evaluated using correlation with human annotators, and thus being inherently subjective.
@@
In this work, we investigate LLM judges on mathematical reasoning datasets. Such tasks require complex multi-step reasoning and judgments can be analyzed through the lense of verifiable solutions, allowing us to investigate the relationship between judge and candidate models in a principled manner. In our setup, LLM Judges are given two answers and they have to classify whether both answers, one of them (which one), or none is correct (see Figure 1. We base our analysis on four large ($>27B$ parameters) LLMs and four small ($<10B$ parameters) LLMs on three mathematical reasoning datasets.
Our experiments contain a detailed performance examination. We find that the best tested judge is LLama 3.1 70B, reaching 60% to 90% judgment performance, depending on the dataset. Our results confirm the intuition that judgment performance is aligned with task difficulty.
We perform a statistical analysis of judgment performance and model quality. We find that the individual task performances of judge and candidate models are highly indicative features of judgment performance, as they explain most of the variance in a linear model (as measured by $R^{2}$). On the subset of questions where the candidate models give one correct and one incorrect answer, we uncover an intriguing correlation between judge performance and candidate models’ task performance, indicating that LLM judges tend to select incorrect answers from better models.
We hypothesize that judges partially base their judgment on linguistic cues rather than solely on the reasoning withinin the answers. We follow literature analysing machine-generated text Shaib et al. (2024) and find that 70%-75% of the judgments can be predicted using simple linguistic features, highlighting the systematicity behind the judge decisions.
Lastly, we analyze practical use cases and discuss usage recommendations. Our experiments suggest that LLM judges reliably detect the model of higher task performance but can not reliably improve task performance. Rather, we find that it is more sensible to use the judge model as an answer generator, and subsequently take the majority vote of all three answers.
In summary, our contributions are as follows:
@@
    We perform an in-depth performance analysis of LLM judges on three diverse mathematical reasoning tasks.
@@
    We identify a correlation between model quality, as measured by task performance, and judgment performance, indicating that LLM judges are biased towards higher-quality models.
@@
    We are able to predict the LLM judgment with 70%-75 accuracy using only stylistic patterns, e.g. N-grams of POS-tags. This indicates that LLMs, to a large degree, judge independently of the reasoning.
@@
    We find that judges reliably detect the model of higher quality but are not able to reliably improve task performance.
@@ 2.1 LLM as Judges
Using LLMs as judges to evaluate text generated by LLMs, including their own outputs, has recently attracted significant interest because it reduces the need for human annotation Zheng et al. (2023). Typically, large state-of-the-art models are used as judges. Applications include the automatic assessment of language model capabilities and, such as ranking models with respect to their competence on a given task Zheng et al. (2023), and reinforcement learning from AI feedback by automatically generating data for preference optimization Bai et al. (2022); Wu et al. (2024).
Various methods exist to make judgments Zheng et al. (2023); Liusie et al. (2024). One approach is pairwise selection Wang et al. (2024b), where two answers are presented, and the model is asked to select the better one. Another approach is pointwise grading Li et al. (2024), where the model is asked to assign a grade based on a predefined scale, and the answer with a better grade is chosen. Judgment prompts may involve reference solutions or not. Another body of research explicitly trains models to act as judges Kim et al. (2024a); Wang et al. (2024b) or closely related, as reward models Wang et al. (2024c); Li et al. (2024).
The effectiveness of LLMs as judges is typically assessed by measuring the correlation or overlap with human judgments Zheng et al. (2023); Kim et al. (2024b). In contrast, we focus on tasks with a concrete final answer. Finally, we want to stress that several works caution against the use of LLM judges as experts Bavaresco et al. (2024); Koo et al. (2023); Raina et al. (2024); Doddapaneni et al. (2024).
@@ 2.2 Biases in LLM-as-a-judge
One linguistic bias is ordering bias Zheng et al. (2023); Koo et al. (2023); Wang et al. (2024a), where a judge gives a different answer depending on the order in which answers are presented. Panickssery et al. (2024) note that it is possible to interpret position bias as a sign that the model is unsure. There are multiple works Xu et al. (2024); Panickssery et al. (2024); Liu et al. (2024) that find evidence for self-bias or self-preference. Koo et al. (2023) provide a benchmark for analyzing cognitive biases. West et al. (2024) and Oh et al. (2024) explore the “Generative AI Paradox” where it is easier for LLMs to generate solutions rather than analyzing them, unlike humans who often find analysis easier than generation.
In this work, we aim to establish a better understanding of underlying patterns that relate judgments to interpretable factors, such as task performance or stylistic patterns.
@@ 3.1 Problem Description
In this work, we use an LLM judge, referred to as $J$, to assess answers produced by two other candidate LLMs, $A$ and $B$, in response to math questions (see Figure 1 for an illustrative example). The two candidate answers may both be correct, both incorrect, or either the answer of model $A$ or $B$ correct. The judge’s task is to determine which of these cases applies by reviewing both the CoT reasoning and the final responses provided in candidate answers.
Thus, the judge engages in a four-class classification task. We denote the judge’s accuracy by the score $S^{J}_{A,B}$ and call this metric judgment performance. Further, we define the task performance of an individual model $X$ on a specific dataset as $S_{X}$, e.g. $S_{A}$, $S_{B}$ or $S_{J}$.
@@ 3.3 Models
We evaluate the performance of openly available LLMs, including four large models including Qwen 2.5 72B Yang et al. (2024), Llama 3.1 70B AI@Meta (2024), Yi 1.5 34B Young et al. (2024), Mixtral 8x7B Jiang et al. (2024) and four small models, namely Llama 3 8B AI@Meta (2024), Gemma 1.1 7B Gemma Team et al. (2024), Mistral 7B v0.3 Jiang et al. (2023), and Mistral 7B v0.1 Jiang et al. (2023). We use the chat- or instruction-tuned model variants and test each model as a candidate answer generator and as a judge. More information is in Appendix A.2.
3.4 Text Generation
This section describes the generation of candidate answers and judgments. Find more information on prompts and hardware details in Appendix A.
@@ 3.4 Text Generation
For each model we sample two CoT solutions using 4-shot prompting by setting the temperature to $0.9$. By generating two answers $a_{1},a_{2}$ from the same model, we can also evaluate judgments of two different answers by the same model.
Llama 3.1 70B Qwen 2.5 72B Qwen 2.5 14B Gemma 2 27B Qwen 2.5 7B Gemma 2 9B Llama 3.1 8B Gemma 2 2B
(1) $S^{J}_{A,B}$ GSM8K 90.05 85.39 89.2 81.96 81.92 83.60 79.96 64.33
AQUA-RAT 74.47 69.26 72.26 68.48 65.09 67.48 66.26 60.97
MATH 61.18 58.03 62.36 55.34 50.70 52.92 50.96 50.35
(2) Same answer GSM8K 95.27 95.46 95.02 92.27 92.86 94.70 88.13 75.50
AQUA-RAT 79.8 77.74 77.19 78.67 77.21 77.94 74.52 76.01
MATH 79.09 77.91 76.86 75.39 73.14 75.65 71.05 77.85
(3) Different Answer GSM8K 70.04 55.67 68.43 47.09 49.93 49.50 51.41 31.71
AQUA-RAT 57.44 48.92 57.64 45.55 40.45 46.31 44.10 30.76
MATH 48.95 45.40 51.47 42.66 36.70 38.86 36.41 32.50
(4) 1-correct GSM8K 78.18 64.08 76.92 52.25 56.19 58.10 59.26 27.78
AQUA-RAT 66.43 57.10 69.22 44.44 47.13 54.43 52.93 24.05
MATH 71.92 70.19 79.62 41.80 57.79 60.97 60.73 22.62
Table 1: Performance of judge LLMs (1) on all samples, (2) on samples where $A$ and $B$ agree, (3) on samples where $A$ and $B$ disagree and (4) on samples where exactly one given answer is correct. Results are averaged over all candidate model pairs $(A,B)$. The highest accuracy is bold and the second highest underlined.
@@
For all $36$ unique model combinations $(A,B)$ [^2] We consider all pairs from the eight LLMs, including self-pairing, yielding $\binom{8+2-1}{2}=36$ combinations., each model as judge $J$ and each sample of a dataset, we generate a zero-shot judgment. In the case of self-pairing, i.e., $A=B$, we use both generated candidate answers, $a_{1}$ and $a_{2}$. Otherwise, for consistency, we always use the same sampled candidate answer $a_{1}$. We accommodate positional bias Zheng et al. (2023); Koo et al. (2023) by prompting in both possible orders. We obtain the judgment performance by averaging how often the judgments were correctly classified across orderings.
4 Performance Analysis
The experiments have multiple degrees of freedom, such as judges, candidate models, and datasets. To gain a comprehensive understanding of judges’ behavior, we consider two perspectives. First, we investigate judge performance for each dataset, aiming to associate judge performance with task difficulty. Second, for a fixed dataset, we analyze judge performance across different pairs of candidate models.
4.1 General Performance
First, we compare how often the judges make a correct classification across different datasets and different subsets of the datasets.
@@ 4.1 General Performance
We analyze multiple cases, each corresponding to a specific subset of the data. Case (1) investigates the observed judgment performance $S^{J}_{A,B}$ on the full dataset and Case (2) analyzes the subset where both models give the same answer ($A=B$). Case (3) shows the performance where both models give a different answer $(A\neq B$) and Case (4) describes the performance on the subset where exactly one answer is correct. The results are shown in Table 1. Further, we show the class confusion matrix for the four best-performing judges in Figure 2.
@@ Results.
In general, we observe in Table 1 that larger models outperform smaller models, with LLama 3.1 70B performing the best. Interestingly, Qwen 2.5 14B outperforms Qwen 2.5 72B. As shown in Figure 2, the LLM judges have a performance of larger than 95% if both answers are correct. Conversely, the most challenging situation is when both answers are incorrect. It seems that the difficulty of a problem also transcends the difficulty of making a judgment. This is not necessarily intuitive. For instance, humans may find it easier to detect individual wrong reasoning steps and identify wrong answers, respectively.
(a) Qwen2.5 72B
(b) Qwen 2.5 14B
(c) Llama 3.1 70B
(d) Gemma 2 27B
Figure 2: Class confusion matrices per model. We observe that it is difficult for judges to detect that both answers are incorrect.
In cases where one answer is correct and one answer is incorrect, we observe a moderate performance of the judges, reaching up to 80% accuracy (see Case (4) in Table 1 and Figure 2).
In Case (3) where both answers disagree, we observe moderate performance for large models of up to 70%. Here, the smallest model Gemma 2 2B, has a low performance of around 35%. In what follows, we mostly focus on the analysis of the four largest LLMs as judges.